OpenStudy (anonymous):

Solve for the variable. a - (-98) = -2

OpenStudy (anonymous):

a=-100

OpenStudy (anonymous):

\[a + 98 = -2 \implies a = -2-98 \implies a = -100\]

OpenStudy (australopithecus):

@powerangers69 Instead of giving the answer why not show how you got the solution a -1(-98) = -2 a negative * a negative = a positive so a + 98 = -2

OpenStudy (australopithecus):

as anyone can type a problem into http://www.wolframalpha.com/ for a solution